how many liters of a35% solution of alcohol should be mixed with 30 liters of 55% solution to make up a 47% solution
Answer by Maths68(1474)

Soluton A
Amount = x
Concentration = 35%=0.35
================================================================================
Solution B
Amount = 30 liters
Concentration = 55%=0.55
================================================================================
Resultant Solution
Amount=30+x
Concentration=47%=0.47
================================================================================
(Amount of A * Concentration of A)+(Amount of B * Concentration of B) = (Amount of Resultant * Concentration of Resultant)
(x*.35)+(30*0.55)=(30+x)*0.47
0.35x-0.47x=14.1-16.5
-0.12x=-2.4
-0.12x/-0.12=-2.4/-0.12
x=20
================================================================================
20 liters of 35% solution of alcohol should be mixed with 30 liters of 55% solution to make up a 50 liters of 47% solution